# Topological potentials guiding protein self-assembly

### v2

@@ Abstract
The simulated self-assembly of molecular building blocks into functional complexes is a key area of study in computational biology and materials science. Self-assembly simulations of proteins, driven by short-range non-polar interactions, can find the biologically correct assembly as the energy minimizing state. Short-ranged potentials produce rugged energy landscapes however, which lead to simulations becoming trapped in non-functional, local minimizers.
Successful self-assembly simulations depend both on the physical realism of the driving potentials as well as their ability to efficiently explore the configuration space.
We introduce a long-range topological potential, quantified via weighted total persistence, and combine it with the morphometric approach to solvation-free energy. This combination improves the assembly success rate in simulations of the tobacco mosaic virus dimer and other protein complexes by up to sixteen-fold compared with the morphometric model alone. It further enables successful simulation in systems that don’t otherwise assemble during the examined timescales.
Compared to previous topology-based work, which has been primarily descriptive, our approach uses topological measures as an active energetic bias that is independent of electrostatics or chemical specificity and depends only on atomic coordinates. Therefore the method can, in principle, be applied to arbitrary systems where such coordinates are optimized. Integrating topological descriptions into an energy function offers a general strategy for overcoming kinetic barriers in molecular simulations, with potential applications in drug design, materials development, and the study of complex self-assembly processes.
@@ 1 Introduction
In silico assembly of molecular building blocks into functional complexes is a central challenge in biology and materials science. A primary obstacle is the problem of kinetic trapping Perlmutter and Hagan (2014); Varela et al. (2019): even when the functional native structure represents the potential energy minimum, the vastness of the configurational landscape, containing many non-functional local minima, can prevent simulations from discovering it in a feasible timescale. A successful simulation methodology must therefore not only be based on an accurate potential energy function that identifies the native state, but also navigate this rugged landscape efficiently to avoid dead-end assembly pathways.
We introduce a long-range potential based on the principles of computational topology. Topological data analysis, particularly persistent homology Edelsbrunner et al. (2002); Zomorodian and Carlsson (2005), is utilized to solve optimization tasks in machine learning contexts Carrière et al. (2021); Nigmetov and Morozov (2024), and has been widely used to analyze static and dynamic molecular data Cang and Wei (2017); Cang et al. (2018); Benjamin et al. (2023); Xia and Wei (2014) and nanomaterials Hiraoka et al. (2016); Buchet et al. (2018) post hoc. We use it here as an active guiding force within the simulation of protein assembly.
The physically motivated core of our simulations is the morphometric approach to solvation-free energy Mecke (1996); König et al. (2004). It is an implicit solvent model Roux and Simonson (1999); Zhang et al. (2017) and a sophisticated implementation, describing non-polar contributions to the solvation-free energy as a linear combination of four geometric measures: the excluded volume $V$, solvent accessible surface area $A$, integrated mean curvature $C$, and integrated Gaussian curvature $X$ of the solvent accessible surface Lee and Richards (1971); Shrake and Rupley (1973); Richards (1977). This method has proven highly accurate for calculating solvation properties of complex biomolecules Harano et al. (2013); Hansen-Goos et al. (2007); Roth et al. (2006); Evans and Roth (2014) and assembles exotic geometries for generic solutes Spirandelli et al. (2024); Coles and Evans (2024). For simulation of macromolecules, we augment it with a soft-sphere overlap penalty $L$, yielding a total short-range geometric potential:
@@
The prefactors for the geometric terms are derived from fundamental measure theory Hansen-Goos and Roth (2006), where $p$ is pressure, $\sigma$ is surface tension, and $\kappa$ and $\overline{\kappa}$ are bending rigidities König et al. (2004). The weight $\phi$ is empirically determined and, in combination with $L$, acts both as a simple model of repulsion and as a relaxation of the shape of the modeled molecules.
Previous work has shown that for tobacco mosaic virus (TMV) subunits, the assembly observed experimentally corresponds to the minimum of $F_{sol}^{*}$ obtained via Random Walk Metropolis simulations Spirandelli and Evans (2025). This indicates the thermodynamic accuracy of the method. However, this does not guarantee kinetic accessibility. The morphometric approach is inherently short-ranged, as its geometric measures are defined by a local solvent probe (e.g., $r_{s}=1.4\text{\AA }$ for water). Consequently, its energy landscape is largely flat until the subunits are in close contact, at which point it becomes rugged. A simulation driven only by $F_{sol}^{*}$ performs a largely random search until protein subunits interact at a close range. Here, the simulations find local minimizers that are sometimes difficult to escape from, preventing further exploration and correct assembly. This highlights a kinetic, rather than thermodynamic, hurdle that the model has to overcome in a simulation context.
The calculation of morphometric measures already relies on weighted Alpha complexes Edelsbrunner and Mücke (1994); Edelsbrunner (1992); Koehl et al. (2023); Koehl (2021), which naturally provide a multi-scale representation of shape known as a filtration (see Fig. 1). This Alpha-filtration is constructed on the union of atom centers of the simulated molecules. Persistent homology is the tool for quantifying topological features within such a filtration.
@@
The potential captures some shape information over arbitrary ranges suggesting a synergistic strategy in which the two potentials play complementary roles. The long-range topological term, $\mathcal{T}$, can create a guiding funnel across the vast, flat energy landscape, pulling subunits into the correct general vicinity and orientation. The short-range physical potential, $F_{sol}^{*}$, then acts as a refiner, selecting the precise low-energy binding interface. In this work, we test this principle by simulating dimer and trimer assembly of various systems with a combined potential energy function.
$E_{comb}=\mu F_{sol}^{*}+(1-\mu)\mathcal{T}$ (3)
We show that this approach, which combines local physical accuracy with global topological guidance, enhances the purely geometric model, increasing the simulation success rate for the TMV and SARS-CoV-2 ORF9b proteins by more than an order of magnitude and further making the successful assembly of a hepatitis B core protein and an extracellular fragment of the human CD40 ligand possible on the simulated timescales. This result validates the use of topology not just as a post-hoc analytical tool, but as an active component within a potential energy function to solve kinetic challenges in molecular simulation.
@@ 2.1 Geometry-Topology Correlations
Figure 2: Comparing potentials of a sequence of states obtained via a simulation driven by the geometric potential $F^{*}_{sol}$. The geometric itself (A). A goodness of fit measure with respect to the correct assembly of two TMV proteins (B). The weighted sum $\mathcal{T}=0.5P_{0}-0.35P_{1}-0.35P_{2}$ (C) The total persistence in dimension zero (D), dimension one (E) and dimension two (F).
We compare the topological and geometric potential, calculating $\mathcal{T}$ for a sequence of configurations of two TMV capsid protein subunits, obtained via a successful simulation run driven by the geometric potential alone. Here successful means that the two subunits are assembled in a configuration similar to one observed experimentally, as the simulation reaches the potential energy minimum. Formally let $S_{ca}$ be an experimentally determined configuration of two correctly assembled protein subunits. Let $S$ be a configuration obtained via a simulation. We say that $S$ and $S_{ca}$ are similar if $d_{\text{RMSD}}(S,S_{ca})<2.5\text{\AA }$. See the Materials and methods section for an explanation of the distance measure $d_{\text{RMSD}}$. A visualization of the correct assembly of two protein subunits is shown in Figure 3 A. Each protein subunit is modeled as a rigid body and its geometry, i.e. the position of its atom centers, is based on the protein data base entry $\operatorname{6R7M}$ Schmidli et al. (2019). This entry also specifies how proteins are positioned in relation to one another in the experimentally observed helical assembly. For atomic radii we take values that subsume the hydrogen atoms into the atoms they are bound to Tsai et al. (1999). The simulation is driven by a Random Walk Metropolis algorithm, see Materials and methods.
@@ 2.1.1 Calculating total persistence for a successful simulation run
Figure 2 A shows how $F_{sol}^{*}$, Equation 1, changes over the course of around $1500$ iterations of a simulated assembly. The signal is flat for large stretches of the simulation because the geometric potential is short-ranged and only exhibits changes in value when the solvent accessible surfaces of the simulated proteins overlap. There are several small valleys in the energy landscape, which correspond to configurations where the subunits are close. They detach again several times before finding a sequence of moves leading to a steep decline in potential energy and a local minimum, corresponding to configurations that are close to experimentally observed assembly. Figure 2 B shows the goodness-of-fit measure $d_{\text{RMSD}}$ that tracks how close to correct assembly the current configuration is. We can see that $d_{\text{RMSD}}$ is more stable in the regions where the subunits overlap and oscillates in the regions where $F_{sol}^{*}$ is flat. Around step $i=1000$ there is a drop in $d_{\text{RMSD}}$ corresponding to the protein subunits being correctly aligned but not yet in contact. As $F_{sol}^{*}$ reaches its minimum, so does $d_{\text{RMSD}}$, which fluctuates around a low value indicating proximity to the correct assembly.
@@
To quantify these observations., we calculate the Pearson coefficient Freedman et al. (2007) $r$ of the different functions. It is a measurement of the linear correlation between two variables $X=\{x_{1},\dots,x_{n}\}$ and $Y=\{y_{1},\dots,y_{n}\}$ and is given by
@@ 2.3 Combining geometry and topology
We will now compare Random Walk Metropolis simulations in which a combination of geometric and topological potential is minimized. We recall the combined potential energy function, Eq. 3:
@@ 2.3.2 Scanning the parameter space for TMV dimerization
Figure 4 C-F shows heat maps for the choices of $(\lambda_{0},\lambda_{1},\lambda_{2})$ mentioned above, where we alter the simulation temperature $T=3,4,5,6$ and the factor $\mu=0.1,0.2,\dots,0.9$ that interpolates between $F_{sol}^{*}$ and $\mathcal{T}$ to calculate the total potential energy as $E_{comb}=\mu F_{sol}^{*}+(1-\mu)\mathcal{T}$. For each point in the heatmap, we run $40$ simulations, where each simulation is run for $50000$ iterations. This corresponded to roughly $12$ hour long simulations on a single core of a high performance cluster.
For the pure morphometric approach, the best success rate we achieve is $5\%$. The highest success rate among mixtures of both the morphometric approach and the topological potential is $85\%$ for $\mathcal{T}_{(1.0,-0.7,-0.7)}$ (Figure 4 D) at a simulation temperature of $5.0K$ and a weighting $\mu=0.5$, as well as for $\mathcal{T}_{(1.0,-0.4,-0.9)}$ (Figure 4 E) at $T=5.0K$ and $\mu=0.3$. This corresponds to a $16$-fold increase in the success rate.
The highest success rate for $\mathcal{T}_{(1.0,-1.0,1.0)}$ (Figure 4 C) is equal to $22.5\%$. It is found for $T=3.0K$ and $\mu=0.8$. Although substantially lower than for the other variants, it is still a $4$-fold increase over the pure morphometric approach.
The combination of $F_{sol}^{*}$ with $\mathcal{T}_{(1.0,0.0,0.0)}$ (Figure 4 F) achieves its best success rate of $12.5\%$ at $T=3.0K$ and $\mu=0.8$. From this we infer that the improvements in success rate are not obtained only because zero-dimensional persistence is pulling the subunits together.
Figure 5: Analysis of the final steps of a simulation driven by $E_{comb}=\frac{5}{10}F_{sol}^{*}+\frac{5}{10}\mathcal{T}_{(1.0,-0.7,-0.7)}$. $F_{sol}^{*}$ as given by Equation 1 (A). $\mathcal{T}_{(1.0,-0.7,-0.7)}$ (B). Images of the configuration at selected states 0, $133$ and $148$ (C-E). The corresponding persistence bar plots with zero-dimensional features in blue, one-dimensional features in green and two-dimensional features in red (F-H).
@@ 2.3.3 Analysis of a successful simulation of the combined potential
Figure 5 illustrates a sample simulation of two subunits with a mixture at $\mu=0.5$ and $T=2.0K$ for $\mathcal{T}=P_{0}-0.7P_{1}-0.7P_{2}$. The simulation reaches its minimum potential energy, a correctly assembled configuration, after $276$ steps. The figures in subfigure A-C show $F_{sol}^{*}$, $\mathcal{T}$ and their mixture relative to the initial (dispersed) state. The D-E subfigures show the individual persistence measures. The D-E subfigures show the persistence barplots for three selected configurations. Namely, the initial state ($i=0$), a state in which the subunits are correctly aligned but have a gap between them ($i=217$) and finally the minimal potential energy state ($i=276$), in which the subunits are correctly assembled. Renders of the corresponding configurations are shown in G-I.
@@
In the final phase, both the topological and geometric energies rapidly decline, while the final gap between the protein subunits is closed. Considering just the topological potential, we observe that the final phase accounts for one third of the total decline, whereas the other two thirds occur during topologically driven alignment of the proteins. Looking at the persistence diagrams in subfigures J-L, we observe that more persistence features appear as the subunits get closer to the correct assembly and that the persistence features are born at lower $\alpha$ values and persist for a shorter time period.

### v3

@@ Abstract
The simulated self-assembly of molecular building blocks into functional complexes is a key area of study in computational biology and materials science. Self-assembly simulations of proteins using physically-motivated potentials for non-polar interactions, can identify the biologically correct assembly as the energy-minimizing state. Short-range potentials, however, produce rugged energy landscapes, which lead to simulations becoming trapped in non-functional local minimizers.
Successful self-assembly simulations depend on the physical realism of the driving potentials as well as their ability to efficiently explore the configuration space.
We introduce a long-range topological potential, quantified via weighted total persistence, and combine it with the morphometric approach to solvation-free energy.
This combination improves the assembly success rate in simulations of the tobacco mosaic virus dimer and other protein complexes by up to sixteen-fold compared with the morphometric model alone. It further enables successful simulation in systems that don’t otherwise assemble during the examined timescales.
Compared to previous topology-based work, which has been primarily descriptive, our approach uses topological measures as an active energetic bias that is independent of electrostatics or chemical specificity and depends only on atomic coordinates. Therefore, the method can, in principle, be applied to arbitrary systems where such coordinates are optimized. Integrating topological descriptions into an energy function offers a general strategy for overcoming kinetic barriers in molecular simulations, with potential applications in drug design, materials development, and the study of complex self-assembly processes.
@@ 1 Introduction
In silico assembly of molecular building blocks into functional complexes is a central challenge in biology and materials science. A primary obstacle is the problem of kinetic trapping [1, 2]: even when the functional native structure represents the minimum of the energy function that governs the simulation, the vastness of the configurational landscape, containing many non-functional local minima, can prevent simulations from discovering it on a feasible timescale. A successful simulation methodology must therefore not only be based on an accurate energy function that identifies the native state, but also navigate it’s landscape efficiently to avoid dead-end assembly pathways.
Here we introduce a long-range potential based on the principles of computational topology to act as a guiding bias. Topological data analysis, particularly persistent homology [3, 4], is utilized to solve optimization tasks in machine learning contexts [5, 6], and has been widely used to analyze static and dynamic molecular data [7, 8, 9, 10] and nanomaterials [11, 12] post hoc. We use it here as an active component within the simulation of protein assembly.
The physically motivated core of our simulations is the morphometric approach to solvation-free energy [13, 14]. It is an implicit solvent model [15, 16] and a sophisticated implementation, describing non-polar contributions to the solvation-free energy as a linear combination of four geometric measures: the excluded volume $V$, solvent accessible surface area $A$, integrated mean curvature $C$, and integrated Gaussian curvature $X$ of the solvent accessible surface [17, 18, 19]. This method has proven highly accurate for calculating solvation properties of complex biomolecules [20, 21, 22, 23] and assembles exotic geometries for generic solutes [24, 25]. For simulation of macromolecules, we augment it with a soft-sphere overlap penalty $L$, yielding a total short-range geometric potential:
@@
The prefactors for the geometric terms are derived from fundamental measure theory [26], where $p$ is pressure, $\sigma$ is surface tension, and $\kappa$ and $\overline{\kappa}$ are bending rigidities [14]. The weight $\phi$ is empirically determined and, in combination with $L$, acts both as a simple model of repulsion and as a relaxation of the shape of the modeled molecules.
Previous work has shown that for tobacco mosaic virus (TMV) subunits, the assembly observed experimentally corresponds to the minimum of $F_{sol}^{*}$ obtained via Random Walk Metropolis simulations [27]. This indicates the thermodynamic accuracy of the method. However, this does not guarantee kinetic accessibility. The morphometric approach is inherently short-ranged, as its geometric measures are defined by a local solvent probe (e.g., $r_{s}=1.4\text{\AA }$ for water). Consequently, its energy landscape is largely flat until the subunits are in close contact, at which point it becomes rugged. A simulation driven only by $F_{sol}^{*}$ performs a largely random search until protein subunits interact at a close range. Here, the simulations find local minimizers that are sometimes difficult to escape from, preventing further exploration and correct assembly. This highlights a kinetic, rather than thermodynamic, hurdle that the model has to overcome in a simulation context.
@@
The calculation of morphometric measures already relies on weighted Alpha complexes [28, 29, 30, 31], which naturally provide a multi-scale representation of shape known as a filtration (see Fig. 1). This Alpha-filtration is constructed on the union of atom centers of the simulated molecules. Persistent homology is the tool for quantifying topological features within such a filtration.
@@
The potential captures some shape information over arbitrary ranges suggesting a synergistic strategy in which the two potentials play complementary roles. The long-range topological term, $\mathcal{T}$, can create a guiding funnel within the energy landscape, pulling subunits into the correct general vicinity and orientation. The short-range geometric potential, $F_{sol}^{*}$, then acts as a refiner, selecting the precise low-energy binding interface. In this work, we test this principle by simulating dimer and trimer assembly of various systems governed by a combined potential:
$E_{comb}=\mu F_{sol}^{*}+(1-\mu)\mathcal{T}.$ (3)
We show that this approach, which combines local physical accuracy with global topological guidance, enhances the purely geometric model, increasing the simulation success rate for the TMV and SARS-CoV-2 ORF9b proteins by more than an order of magnitude and further making the successful assembly of a hepatitis B core protein and an extracellular fragment of the human CD40 ligand possible on the simulated timescales. This result validates the use of topology not just as a post-hoc analytical tool, but as an active biasing potential to solve kinetic challenges in molecular simulation.
@@ 2.1 Geometry-Topology Correlations
Figure 2: Comparing values of a sequence of states obtained via a simulation driven by the geometric potential $F^{*}_{sol}$. The geometric potential itself (A). A goodness of fit measure with respect to the correct assembly of two TMV proteins (B). The weighted sum $\mathcal{T}=0.5P_{0}-0.35P_{1}-0.35P_{2}$ (C) The total persistence in dimension zero (D), dimension one (E) and dimension two (F).
We compare the topological and geometric potential, calculating $\mathcal{T}$ for a sequence of configurations of two TMV capsid protein subunits, obtained via a successful simulation run driven by the geometric potential alone. Here successful means that the two subunits are assembled in a configuration similar to one observed experimentally, as the simulation reaches the energy minimum. Formally let $S_{ca}$ be an experimentally determined configuration of two correctly assembled protein subunits. Let $S$ be a configuration obtained via a simulation. We say that $S$ and $S_{ca}$ are similar if $d_{\text{RMSD}}(S,S_{ca})<2.5\text{\AA }$. See the Materials and methods section for an explanation of the distance measure $d_{\text{RMSD}}$. A visualization of the correct assembly of two protein subunits is shown in Figure 3 A. Each protein subunit is modeled as a rigid body and its geometry, i.e. the position of its atom centers, is based on the protein data base entry $\operatorname{6R7M}$ [32]. This entry also specifies how proteins are positioned in relation to one another in the experimentally observed helical assembly. For atomic radii we take values that subsume the hydrogen atoms into the atoms they are bound to [33]. The simulation is driven by a Random Walk Metropolis algorithm, see Materials and methods.
@@ 2.1.1 Calculating total persistence for a successful simulation run
Figure 2 A shows how $F_{sol}^{*}$, Equation 1, changes over the course of around $1500$ steps of a simulated assembly pathway. Here, a step of the assembly pathway corresponds to an accepted proposal of the Random Walk Metropolis algorithm, corresponding to a change in configuration space. The signal is flat for large stretches of the simulation because the geometric potential is short-ranged and only exhibits changes in value when the solvent accessible surfaces of the simulated proteins overlap. There are several small valleys in the energy landscape, which correspond to configurations where the subunits are close. They detach again several times before finding a sequence of moves leading to a steep decline in energy and a local minimum, corresponding to configurations that are close to experimentally observed assembly. Figure 2 B shows the goodness-of-fit measure $d_{\text{RMSD}}$ that tracks how close to correct assembly the current configuration is. We can see that $d_{\text{RMSD}}$ is more stable in the regions where the subunits overlap and oscillates in the regions where $F_{sol}^{*}$ is flat. Around step $i=1000$ there is a drop in $d_{\text{RMSD}}$ corresponding to the protein subunits being correctly aligned but not yet in contact. As $F_{sol}^{*}$ reaches its minimum, so does $d_{\text{RMSD}}$, which fluctuates around a low value indicating proximity to the correct assembly.
@@
To quantify these observations, we calculate the Pearson coefficient [34] $r$ of the different functions. It is a measurement of the linear correlation between two variables $X=\{x_{1},\dots,x_{n}\}$ and $Y=\{y_{1},\dots,y_{n}\}$ and is given by
@@ 2.3 Combining geometry and topology
We will now compare Random Walk Metropolis simulations in which a combination of geometric and topological potential is minimized. We recall the combined potential, Eq. 3:
@@ 2.3.2 Scanning the parameter space for TMV dimerization
Figure 4 C-F shows heat maps for the choices of $(\lambda_{0},\lambda_{1},\lambda_{2})$ mentioned above, where we alter the simulation temperature $T=3,4,5,6$ and the factor $\mu=0.1,0.2,\dots,0.9$ that interpolates between $F_{sol}^{*}$ and $\mathcal{T}$ to calculate the total energy as $E_{comb}=\mu F_{sol}^{*}+(1-\mu)\mathcal{T}$. For each point in the heatmap, we run $40$ simulations, where each simulation is run for $50000$ iterations. This corresponded to roughly $12$ hour long simulations on a single core of a high performance cluster.
For the pure morphometric approach, the best success rate we achieve is $5\%$. The highest success rate among mixtures of both the morphometric approach and the topological potential is $85\%$ for $\mathcal{T}_{(1.0,-0.7,-0.7)}$ (Figure 4 D) at a simulation temperature of $5.0$ and a weighting $\mu=0.5$, as well as for $\mathcal{T}_{(1.0,-0.4,-0.9)}$ (Figure 4 E) at $T=5.0$ and $\mu=0.3$. This corresponds to a $16$-fold increase in the success rate.
The highest success rate for $\mathcal{T}_{(1.0,-1.0,1.0)}$ (Figure 4 C) is equal to $22.5\%$. It is found for $T=3.0$ and $\mu=0.8$. Although substantially lower than for the other variants, it is still a $4$-fold increase over the pure morphometric approach.
The combination of $F_{sol}^{*}$ with $\mathcal{T}_{(1.0,0.0,0.0)}$ (Figure 4 F) achieves its best success rate of $12.5\%$ at $T=3.0$ and $\mu=0.8$. From this we infer that the improvements in success rate are not obtained only because zero-dimensional persistence is pulling the subunits together.
Figure 5: Analysis of the final steps of a simulation driven by $E_{comb}=\frac{5}{10}F_{sol}^{*}+\frac{5}{10}\mathcal{T}_{(1.0,-0.7,-0.7)}$. All energy plots (A-F) show values relative to step 0. We see $F_{sol}^{*}$ as given by Equation 1 (A), $\mathcal{T}_{(1.0,-0.7,-0.7)}$ (B), $E_{comb}$ (C), and the individual persistence measures (D-F). We show persistence bar plots of steps 0, $217$ and $276$ with zero-dimensional features in blue, one-dimensional features in green and two-dimensional features in red (J-L). The corresponding protein configurations are shown in (G-I).
@@ 2.3.3 Analysis of a successful simulation of the combined potential
Figure 5 illustrates a sample simulated assembly pathway of two subunits with a mixture at $\mu=0.5$ and $T=2.0$ for $\mathcal{T}=P_{0}-0.7P_{1}-0.7P_{2}$. Again, steps of the assembly pathway refer to accepted proposals of the Random Walk Metropolis algorithm, corresponding to a change in configuration space. The simulation reaches its minimum energy, a correctly assembled configuration, after $276$ steps. The figures in subfigure A-C show $F_{sol}^{*}$, $\mathcal{T}$ and their mixture relative to the initial (dispersed) state. The D-E subfigures show the individual persistence measures. The D-E subfigures show the persistence barplots for three selected configurations. Namely, the initial state ($i=0$), a state in which the subunits are correctly aligned but have a gap between them ($i=217$) and finally the minimal energy state ($i=276$), in which the subunits are correctly assembled. Renders of the corresponding configurations are shown in G-I.
@@
In the final phase, both the topological and geometric potentials rapidly decline, while the final gap between the protein subunits is closed. Considering just the topological potential, we observe that the final phase accounts for one third of the total decline, whereas the other two thirds occur during topologically driven alignment of the proteins. Looking at the persistence diagrams in subfigures J-L, we observe that more persistence features appear as the subunits get closer to the correct assembly and that the persistence features are born at lower $\alpha$ values and persist for a shorter time period.